OpenStudy (monkey*):

Given f(x)=-x^2+2x, find f(-3).

OpenStudy (monkey*):

Help?

Nnesha (nnesha):

f(-3) means x=-3 so substitute all x's for -3

OpenStudy (monkey*):

So, f(-3)=-3^2+2(-4) ?

OpenStudy (monkey*):

(-3)**

Nnesha (nnesha):

yes right remember coefficient is also negative so \[\huge\rm -(-3)^2+2(-3)\] now solve

OpenStudy (monkey*):

3?

Nnesha (nnesha):

hmm no (-3)^2 =?

OpenStudy (monkey*):

-6?

Nnesha (nnesha):

(-3)^2 is same as -3 times -3

OpenStudy (monkey*):

9.

Nnesha (nnesha):

yes right so-(9)+2(-3) = ?

OpenStudy (monkey*):

I got -15.

Nnesha (nnesha):

right

OpenStudy (monkey*):

Would that be my final answer?

Nnesha (nnesha):

yes

OpenStudy (monkey*):

Okay, would you mind helping me with just one more?

Nnesha (nnesha):

i'll try

OpenStudy (monkey*):

If f(x)=x-3 and g(x)=2x-4, find (f+g)(x).

OpenStudy (monkey*):

I tried working this one out and got 3x-7. Which I don't think is right.

Nnesha (nnesha):

just one question how did you get 3x :=)

OpenStudy (monkey*):

I have no idea honestly. Haha

Nnesha (nnesha):

so how did you get 3x-7 ?

OpenStudy (monkey*):

At the end how it has f+g and then x. I think that's how I got the x. Can you help?

Nnesha (nnesha):

(f+g)(x) means add both function f(x) and g(x) so f(x) + g(x)

Nnesha (nnesha):

so replace f(X) with x-3 and f(x) with 2x-4 and then combine like terms (x-3)+(2x-4)

OpenStudy (monkey*):

Wouldn't you add the x to 2x and it becomes 3x?

Nnesha (nnesha):

yes right that's what i was asking

Nnesha (nnesha):

good job!

OpenStudy (monkey*):

Thank you, so now the -3 and -4¿

OpenStudy (monkey*):

?

Nnesha (nnesha):

yes right -3-4 are like terms

OpenStudy (monkey*):

Would it be -7?

OpenStudy (monkey*):

Ohh, would my answer actually be 3x-7?

Nnesha (nnesha):

yes right

OpenStudy (monkey*):

Thank you so much for the help!

Nnesha (nnesha):

yw :=)